OpenStudy (anonymous):

cos(θ) = 33/65 where 0 < θ < π/2 sin θ/2= cos θ/2=

OpenStudy (jdoe0001):

same as before http://openstudy.com/study#/updates/528d7493e4b0fd0888879dc5

OpenStudy (jdoe0001):

you're in the 1st Quadrant this time, thus sine and cosine will both be positive use the pythagorean theorem to find side "b" once you have "a", "b" and "c", use the half-angle identities for sine and cosine and plug in your sine and cosine values for \(\theta\)

OpenStudy (anonymous):

yeah, i know what to do but my answer isn't right

OpenStudy (jdoe0001):

hnmmm

OpenStudy (anonymous):

yeah

OpenStudy (jdoe0001):

actually, I just recall, you don't even need to do all that, since both half-angle identities just use cosine....

OpenStudy (anonymous):

yeah, i used 33/55 and still didn't get the right answer

OpenStudy (jdoe0001):

one sec

OpenStudy (anonymous):

okay

OpenStudy (jdoe0001):

\(\bf cos(\theta)=\cfrac{33}{65}\qquad \qquad \pi<\theta < \frac{\pi}{2}\\ \quad \\ sin\left(\frac{\theta}{2}\right)=\sqrt{\cfrac{1-cos(\theta)}{2}}\implies sin\left(\frac{\theta}{2}\right)=\sqrt{\cfrac{1-\frac{33}{65}}{2}}\\ \quad \\ sin\left(\frac{\theta}{2}\right)=\sqrt{\cfrac{\frac{32}{65}}{2}}\implies sin\left(\frac{\theta}{2}\right)=\sqrt{\cfrac{16}{65}}\implies sin\left(\frac{\theta}{2}\right)=\cfrac{\sqrt{16}}{\sqrt{65}}\\ \quad \\ sin\left(\frac{\theta}{2}\right)=\cfrac{4}{\sqrt{65}}\implies sin\left(\frac{\theta}{2}\right)=\cfrac{4}{\sqrt{65}}\cdot \cfrac{\sqrt{65}}{\sqrt{65}}\implies sin\left(\frac{\theta}{2}\right)=\cfrac{4\sqrt{65}}{65}\)

OpenStudy (jdoe0001):

\(\bf cos(\theta)=\cfrac{33}{65}\qquad \qquad \pi<\theta < \frac{\pi}{2}\\ \quad \\ \quad \\ cos\left(\frac{\theta}{2}\right)=\sqrt{\cfrac{1+cos(\theta)}{2}}\implies cos\left(\frac{\theta}{2}\right)=\sqrt{\cfrac{1+\frac{33}{65}}{2}}\\ \quad \\ cos\left(\frac{\theta}{2}\right)=\sqrt{\cfrac{\frac{98}{65}}{2}}\implies cos\left(\frac{\theta}{2}\right)=\sqrt{\cfrac{49}{65}}\implies cos\left(\frac{\theta}{2}\right)=\cfrac{\sqrt{49}}{\sqrt{65}}\\ \quad \\ cos\left(\frac{\theta}{2}\right)=\cfrac{7}{\sqrt{65}}\implies cos\left(\frac{\theta}{2}\right)=\cfrac{7}{\sqrt{65}}\cdot \cfrac{\sqrt{65}}{\sqrt{65}}\implies cos\left(\frac{\theta}{2}\right)=\cfrac{7\sqrt{65}}{65}\)

OpenStudy (anonymous):

thanks

OpenStudy (jdoe0001):

yw